# While recognizing actions in videos, LMMs struggle to detect core interaction events

### v1

@@ 5 Results
S-Table 3: Grounding effect on LMMs’ performance. Prior to the main interaction detection task, models are instructed to name the target object and specify the length of the video sequence, to improve their perceptual grounding. Mean accuracy is measured for the detection of the frame where the test event occurs, within an allowed error tolerance (here the exact true frame or up to 2 frames off). Results are reported as per 3 ICL regimes (ZS, OS, TS) for the models: (i) Qwen-2.5VL-72B and (ii) GPT-4o.
ICL Grounding Mean Accuracy (%)
Exact 2-off
Qwen GPT Qwen GPT
[…]
W $8\pm 2$ $8\pm 2$ $52\pm 4$ $51\pm 4$
OS W/O $9\pm 3$ $12\pm 1$ $\textbf{59}\pm 3$ $63\pm 3$
W $8\pm 1$ $11\pm 1$ $56\pm 3$ $62\pm 3$
TS W/O $13\pm 2$ $14\pm 1$ $63\pm 3$ $67\pm 3$
W $12\pm 2$ $14\pm 1$ $3\pm 3 66$$6\pm 3$
(a) Example 1 - False prediction
(b) Example 2 - Correct prediction
S-Figure 4: Example predictions of the model Qwen-2.5VL-72B. The model provides the presented chain-of-thought under the ”WITH” Reasoning condition. (a) A false prediction. (b) A Correct prediction. The examples show that the reasoning seems logical and realistic, but the relation to the actual video frames is often very loose. Orange and green boxes mark the true frame. Red box marks a false prediction.
Action recognition.
As a comparison with models ability to detect core interaction events, here we tested the models’ ability to recognize the action in the test videos. The SSv2 dataset, includes action template labels, for example, ”Putting something, something and something on the table”, each correspond to many videos in the dataset. In addition, the dataset provides a unique label for each video, where the ”something” placeholder is replaced with the corresponding object names in that video, for example, ”Putting a scale, an eraser and a sd card on the table”.
In this experiment, we evaluated the models on the experimental dataset, which is a subset of the SSv2 dataset, consisting of 33 videos spanning 15 action templates. For template recognition, the prompt included a verbal description of four representative visual examples for a few templates, used solely to illustrate the interaction style. No actual video frames were included. The models were first instructed to classify the general actions depicted in the videos using the predefined list of SSv2 action templates. Importantly, the models were required to output only template-IDs rather than the textual template labels. This allowed us to evaluate predictions by comparing numerical IDs instead of free-form strings, thereby avoiding mismatches caused by formatting, spacing, or minor textual variations.
Using this ID-based evaluation, the Top-1 classification accuracy for the Qwen model was 61 ± 9%, and 48 ± 9% for the ChatGPT model. The Top-5 accuracy was 85 ± 6% and 82 ± 7% for Qwen and ChatGPT, respectively.
In addition, we evaluated whether the models could correctly identify the objects involved in each interaction. For this task, the model was given the correct action template as input and was asked to replace every occurrence of the placeholder “something” with the corresponding object(s) appearing in the video. The number of required object names matched the number of placeholder occurrences in the template. To compute accuracy, the predicted placeholders were compared manually against the ground-truth object labels, allowing for differences in ordering and for reasonable synonyms referring to the same object.
The accuracy of correctly naming all objects in a template was 70 ± 8% for Qwen-2.5VL-72B and 82 ± 7% for GPT-4o.
Event bounding-box detection.
@@
We evaluated the models ability to locate the interaction region within the predicted video frame, where the event occurs. Since the annotations contain a point image location per event (point of contact or release), we considered, for the evaluation, a square bounding box centered around the annotation point. We tested several box sizes in the range $20\times 20$ to $200\times 200$ pixels. We computed the Intersection-over-Union (IoU) between the models’ predicted bounding boxes and the boxes around the true event position. Using a $120\times 120$ pixel size boxes around the true location point yielded mean IoU of only $\textbf{6.2}\%$ – the best performance in the range of boxes used around the true location. None of the predicted bounding boxes yielded IoU above the common 50% detection threshold. Despite the low IoU scores, the true event location points ware included inside the models’ predicted bounding boxes in more than $90\%$ of the cases. However, further inspection revealed that the models often predicted over-sized bounding boxes, which in many cases even exceeded the actual image dimensions, indicating that the models internally rescaled the input frames or ignored the original spatial scale. It should be noted, that we explicitly instructed the model to use the original image scale and respond with the smallest possible bounding box. Overall, these findings indicate that the model is unable to reliably localize interaction regions and tends to output bounding boxes whose spatial scale does not correspond to the input frame.
@@ 6 Discussion
Recent advancement of current computational vision models improve significantly LMMs ability to recognize actions in real-world video sequences with high level of generalization to unseen actions and scenes [tang_llm_video_survey2025, buschoff_nature2025, tong2022videomae]. Our experiments on action recognition verify this increased performance for the models Qwen-2.5VL-72B and GPT-4o (as well as others, see Supplementary). However, previous studies have indicated a possible limitation of current vision models in understanding the interactions and core events underlying the general action [Gunawardhana2024SelfSupervised]. Do LLMs overcome this limitation by leveraging their vast common knowledge and visual semantics?
@@
In this paper we conducted a series of experiments, under several in-cotext-learning (ICL) regimes, to test LMMs’ (Qwen-2.5VL-72B and GPT-4o) ability in detecting where and when in the video, core interaction events occur. Specifically, we focused on ”contact” and ”release” events, where a target object becomes attached to an agent (e.g., a hand) or detached from the agent (see Sec. 4). We introduce a new large scale dataset with more than 20K human annotated events in videos from the SSv2 dataset [goyal2017something] (see Sec. 3).
@@
Our experimental results indicate that despite the ability to classify correctly the action in the videos ($\textbf{80}\%$ in Top-5) and even name the correct target objects in the scenes ($\textbf{70}\%$), the models struggle with detecting the core interaction events and ground them visually by associating a particular frame and image location, where they occur ($\leq\textbf{11}\%$). Introducing similar examples using few-shot ICL paradigm slightly improves the performance, which still remains very low ($\leq\textbf{12}\%$ in OS; $\leq\textbf{15}\%$ with TS).
@@
As demonstrated in earlier studies [wei_cot_prompting2022], applying Chain-of-Thought prompting seems to increase the models’ reasoning ability (”WITH” Reasoning condition in our experiments), but yielded a minor performance increase only for Qwen ($4\%$, see Tab. 2). Interestingly, explicitly instructing the models to attend and describe information in the instructing prompt and the input video (e.g., name the target objects in the interaction scene), did not improve the models’ performance and even yielded lower accuracy in the majority of the experiments (under ”WITH” Grounding condition, see Tab. 3). These results seem to be somewhat in contrast with previous studies such as by Shao et al. [shao_visual_cot2024]. They suggest that this level of visual grounding is not enough to enhance the models’ understanding of core interaction events.
@@
We find that models struggle with the perceptual grounding of the core events underlying actions and interactions in the visual input, despite their general ability to describe the action and participating objects and agents in the interactions. This limitation is partly related to the challenge of complex question decomposition as was already shown in previous studies [zhang_vqa_dec2024]. However, it seems that there is more to this limitation. We hypothesize that the main limitation is rooted in a loose integration between the visual representation (often of pretrained visual transformers) and the language representation, which are mostly trained separately. This limitation projects also to the models’ inability to overcome current challenges of visual models in interpreting spatial relations between objects [liu-etal-2025-multimodal-large] and complex dynamic events, despite their huge semantic knowledge. In a sense, the models are merely able to provide a convincing ”story” about the interactions based on the reliably detected objects in the scenes and the models’ familiarity with verbal action descriptions. In struggling to pinpoint the moment and location of physical contact that defines the interactions, the models lack the perceptual grounding required for deeper understanding of dynamic scenes
@@
The implication of this limitation may be that current LMMs lack the capacity to develop full visual understanding of dynamic interactions, similar to intelligent beings [michotte1963perception, buschoff_nature2025], and thus can have only limited ability in interpreting unfamiliar and complex interactions, as well as in planning interactions on their own for artificial systems.
@@ 7 Conclusions
In this paper we demonstrate a major limitation of current large multi-modal models in understanding dynamic interactions. Despite the ability to describe the action and name the participating objects, the models struggle to point back to the spatiotemporal location of core interaction events in videos. We introduce an extension to the SSv2 dataset with more than $20K$ detailed annotations of such core events in more than $10K$ videos. These annotations may be used in future efforts to develop new foundation models with better understanding of visual dynamic interactions.
[…]
@@ 8.1 Action and object recognition
This section provides extended descriptions of the experimental setup of the evaluation on the tasks related to action and object recognition, mentioned in Section 4 of the main manuscript. All experiments were conducted using the extracted video frames of the full length videos from the original SSv2 dataset [goyal2017something]. The models were given, as input, the sequence of video frames and prompted with the task-specific textual information detailed below. Both Qwen-2.5-VL-72B and GPT-4o were evaluated under the zero-shot regime (without any examples), and without any of the modifying conditions (Grounding and Reasoning).
@@
LABEL:supplst:placeholder_prompt and LABEL:supplst:act_recog_prompt presents the prompts used in the experiments for testing the models performance in the tasks of object and action recognition, respectively. In the listed prompts, the strings template_sentence, template_id, and n_slots denote variables that were automatically replaced during execution for each experiment, according to the specific interaction template. The prompt structure itself was fixed, while these fields changed to reflect the corresponding template text, its numeric ID, and the number of placeholder slots.
@@ 8.2 Tow-Shot with Feedback
Inspired by recent work on boosting LLMs reasoning with verification feedback [wei_cot_prompting2022, yao_tot2023, guo_deepseekr12025], we designed a variant to the common TS regime, in which the label of the second example was presented to the model indirectly, through an iterative feedback session. In this session, the model performed the detection task on the second example video and then prompted a numerical verification feedback by the user side. The feedback indicate a metric on the gap between the predicted frame ($f_{p}$) and the true frame ($f_{t}$) of the event. We defined the error in prediction, $\epsilon(f_{t},f_{p})$, via a sigmoid function, as shown in Eq. 1. The error function is shifted to the middle of the frame range ($\frac{N}{2}$) (as we constrained the error to be in the range of $[0,1]$, and required it to be $\epsilon(\frac{N}{2})=0.5$). The score was defined as shown in Eq. 2. The iterative feedback session ended when the model predicted the correct frame, or after it exceeded a limit of allowed trials $T_{th}$, which in our experiments was set to 10 (equivalent to the maximal number of iterations in the naive case where the model simply scans all the frames in turn until it gets to the right frame). After the iterative session ended, the model was instructed to perform the main detection task on the test video. In our experiment, we included examples only of other events from the same full video from which the test event was cropped, thus providing the model context with familiar context from the test video. The protocol followed the algorithm in Algorithm 1. The instructing prompt is presented in LABEL:lst:ts_fb_prompt.
@@
Listing 5: Instruction prompt for object detection.
@@
Listing 6: Instruction prompt for event bounding-box detection.
@@
Listing 7: Instruction prompt for TS with feedback experiment.
User:
In the next procedure follow these rules:
6) You should never predict the same frame twice.
7) Stop your predictions only when the score is above 0.98.
Listing 8: Instruction prompt for action recognition.

### v2

@@ 5 Results
Table 3: Grounding effect on LMMs’ performance. Prior to the frame detection task, models are instructed to name the target object, to improve their perceptual grounding. Mean accuracy is measured for the detection of the frame where the test event occurs, within an allowed error tolerance (here the exact or one frame off the true frame). Results are reported as per 3 ICL regimes (ZS, OS, TS) for the models: (i) Gemini-2.5-Flash and (ii) GPT-5.2.
ICL GRND Mean Accuracy Percentage (SD)
Exact 1-off
Gemini GPT Gemini GPT
ZS W/O $7.58$ $(1.89)$ 11.62 $(2.28)$ $29.29$ $(3.24)$ 34.85 $(3.39)$
W $9.60$ $(2.10)$ $10.10$ $(2.15)$ $31.31$ $(3.30)$ $34.34$ $(3.38)$
OS W/O $17.78$ $(3.84)$ $17.31$ $(2.18)$ $44.77$ $(3.53)$ $41.94$ $(3.16)$
W 18.84 $(2.78)$ $16.77$ $(2.14)$ 46.09 $(3.54)$ $40.45$ $(3.09)$
TS W/O $19.32$ $(2.81)$ $19.31$ $(2.81)$ 49.49 $(3.55)$ $45.22$ $(3.54)$
W 19.37 $(2.81)$ $15.84$ $(2.60)$ $47.10$ $(3.55)$ $39.92$ $(3.48)$
(a) Mean Accuracy (ZS)
(b) Mean Frame Error (ZS)
(c) GPT-5.2
Figure 5: Models’ performance. (a) A comparison between models’ performance at Zero-Shot ICL. Exact detection of the physical event is around chance level for most models. At one-frame off error tolerance Gemini-pro-3 peeks at 42.9%. Humans outperform the models by far. (b) The models’ mean frame error. A positive error indicates a prediction which precedes a physical contact event, or follows a physical release event. In these frames there is often a clear gap between the hand and the target object. (c) Detail mean accuracy vs. frame error tolerance for GPT-5.2. The chart depicts a slight performance increase across the different ICL regimes (ZS, OS, TS). Results are under the “WITH” Reasoning condition. Chance level is 10% (shown as a dashed horizontal line) as the length of all videos in the experimental dataset is 10 frames.
Fig. 5 presents performance charts for the evaluated models. The results in Fig. 5a indicate that at the ZS regime, the models’ accuracy is about chance level (10%) for detecting the exact frame where the event occurs. The accuracy for an error tolerance of one-frame off peaks at 42.9% with Gemini-3-Pro. However, as indicated by Fig. 5b, the models’ failure often tends to frames with no visual grounding of a physical contact. In addition, when evaluated on ’non-event’ control videos, which contained hands and objects but did not present a physical contact, most of the models completely failed. Except for Qwen, which responded correctly at 18.2% of the trials that there is no event in these video, the rest of the models always predicted a frame and provided an explanation for their choice. The chart in Fig. 5c depicts the GPT-5.2 model’s performance as a function of the allowed error tolerance (other models yielded similar results, see Supplementary). The chart also shows a minor increase in performance between the different ICL regimes as was demonstrated already in earlier studies [8].
Tab. 2 reports the mean accuracy for $e_{\tau}=0$ and $e_{\tau}=1$ as per each of the three ICL regimes (ZS, OS, TS) under the Reasoning condition (see Sec. 4). Similarly, Tab. 3 lists the mean accuracy under the Grounding condition (see supplementary for the results on other models).
These results indicate that explicitly instructing the models to provide visual grounding evidence and reasoning for their answers does not necessarily increase their performance. Surprisingly, the results indicate that these instructions, may sometimes yield a decrease in the models’ performance. This reflects the foundational gap between System 1 and System 2 abilities of the models. We further discuss this in Sec. 6.
Finally, we performed an ablation experiment, inspired by recent work on boosting LLMs reasoning with a verification feedback procedure – similar to reinforcement learning approach [32, 35, 12] – but this did not improve the models’ performance on the detection task (see supplementary).
Action and object recognition.
To contextualize the failure in fine-grained grounding, we benchmarked the models’ high-level action and object recognition capabilities. Using our experimental dataset consisting of a subset of 33 videos across 15 action templates from SSv2, we evaluated two distinct tasks:
(i) Action Classification: Models were prompted to identify the correct SSv2 action template for each video. To ensure evaluation precision and avoid linguistic ambiguity, we required models to output numerical template IDs rather than text strings. Accuracy was measured using Top-1 and Top-5 metrics. For few-shot guidance, the prompt included verbal descriptions of four representative interaction styles, excluding actual video frames.
(ii) Object Naming: Given the correct action template, models were tasked with replacing ”something” placeholders with the specific objects identified in the video. Accuracy was determined by manually verifying predicted objects against ground-truth labels, accounting for ordering and synonyms.
Action Classification Acc. (%) Obj. Naming Acc. (%)
Model Top-1 Top-5
GPT-4o $48.48\pm 8.69$ $81.82\pm 6.71$ $81.82\pm 6.71$
GPT-5.2 $45.45\pm 8.67$ $\textbf{96.97}\pm 2.98$ $93.94\pm 4.15$
Gemini-2.5 $\textbf{66.67}\pm 8.33$ $90.91\pm 5.08$ $93.94\pm 4.15$
Gemini-3-Pro $63.64\pm 8.37$ $87.88\pm 5.68$ $\textbf{96.97}\pm 2.98$
Qwen-2.5VL $60.61\pm 8.51$ $84.85\pm 6.24$ $69.70\pm 8.00$
Table 4: Per-model accuracy statistics for global action classification and naming the correct target objects in each test video.
As shown in Tab. 4, the models demonstrate robust global understanding, with most achieving over 85% Top-5 action accuracy and high object naming precision. This confirms that their struggle to pinpoint interaction events is not due to a lack of semantic or object-level recognition, but rather a specific deficit in spatiotemporal grounding.
Event bounding-box (BBox) detection.
@@
We evaluated the models ability to locate the interaction region within the predicted video frame, where the event occurs. Since the annotations contain a point image location per event (point of contact or release), we considered for the evaluation several formulations for the prediction of the event location, including exact point location and different BBox formats. We report here the best performing setting, which was predicting a BBox, and measuring the Intersection-over-Union (IoU) between the prediction and a square BBox centered around the label point location. We tested several box sizes in the range 20 to 200 pixels. Using a $\times120$ pixel size boxes around the annotation point yielded the best performance: mean IoU of only $0.147$ for Gemini-3-Pro and $0.086$ for GPT-5.2 (see supplementary). Overall, these findings indicate that the model is unable to reliably localize the regions around the physical interaction.
Human performance.
As a comparison to the models performance, we asked two naive humans to perform the same tasks on the experimental dataset. For humans, the mean accuracy on the event frame detection task was $43.00\%\pm 5.04\%$ for the exact frame, and $81.36\%\pm 4.18\%$ for 1-frame off - high above the best performing model as shown in Fig. 5a. The ICC between the human subjects was $0.87$. The accuracy on the ’non-event’ control set was $70.26\%\pm 6.93\%$. The mean IoU for the event location detection was $0.71\pm 0.02$.
@@ 6 Discussion
Recent advancement of current vision models improve significantly LMMs generalization ability in recognizing unseen actions and scenes in real-world videos [28, 24, 30]. Our action recognition evaluation verify this increased performance for five models: Qwen-2.5VL-72B, GPT-4o, GPT-5.2, Gemini-2.5-Flash and Gemini-3-Pro (see Tab. 4). However, previous studies have indicated a possible limitation of current vision models in understanding the core interaction events underlying the general action [11]. Do LMMs overcome this limitation by leveraging their vast common knowledge and visual semantics?
@@
In this study we conducted a series of experiments, under several ICL regimes, to test LMMs’ ability in detecting where and when in the video, core interaction events occur. Specifically, we focused on ‘contact’ and ‘release’ events, where a target object becomes attached to an agent (e.g., a hand) or detached from the agent (Fig. 1; Sec. 4). We introduced the Contact-Release Interaction Dataset (CRID) – a new large scale dataset with more than 20K human annotated events in videos from the SSv2 dataset [10] (Sec. 3).
@@
Our experimental results indicate that despite the System 1 ability of the models to classify correctly the action in the videos ($80\%$ in Top-5) and even name the correct target objects in the scenes ($70\%$), the models struggle with detecting the physical core events and ground them visually in the videos ($17\%$). Introducing similar examples using few-shot ICL paradigm slightly improves the performance, which still remains slightly above chance level (Fig. 5c).
@@
In contrast with earlier studies [32, 25], applying Chain-of-Thought prompting does not necessarily increase the models’ performance (see Tab. 2). Similarly, explicitly instructing the models to attend and describe the input video (e.g., name the target objects in the interaction scene), does not improve the models’ grounding ability and hence the performance (see Tab. 3). These results suggest that the models lack System 2 understanding abilities of core interaction events.
@@
We find that models struggle with the perceptual grounding of the core events underlying actions and interactions in the visual input, despite their general ability to describe the action and participating objects and agents in the interactions. This limitation is partly related to the challenge of complex question decomposition as was already shown in previous studies [37]. However, it seems that there is more to this limitation. We hypothesize that the main limitation is rooted in a loose integration between the visual representation (often of pretrained visual transformers) and the language representation, which are mostly trained separately. This limitation projects also to the models’ inability to overcome current challenges of visual models in interpreting spatial relations between objects [15] and complex dynamic events, despite their huge semantic knowledge. In a sense, the models exhibit a “shortcut learning” behavior and merely able to tell a “good story” about possible interactions when hands appear in proximity of objects in scenes. In struggling to pinpoint the moment and location of physical contact that defines the interactions, the models lack the perceptual grounding required for deeper understanding of dynamic scenes
@@
The implication of this limitation may be that current LMMs lack the capacity to develop full visual understanding of dynamic interactions, similar to intelligent beings [18, 24], and thus can have only limited ability in interpreting unfamiliar and complex interactions, as well as in planning interactions on their own for artificial systems.
@@ 7 Conclusions
In this paper we demonstrate a major limitation of current large multi-modal models in understanding dynamic interactions. Our analysis suggests that current models are operating as sophisticated System 1 engines. They recognize a ‘picking up’ action by the presence of a hand and a cup, but they do not perform the System 2 ‘mental simulation’ required to pinpoint the exact moment of physical attachment. To move toward genuine multi-modal intelligence, future architectures must incorporate structured priors or causal world models that treat interaction events not just as pixels, but as discrete physical state changes (e.g., attend to motion and motion boundaries around the hand and the object). We introduce CRID - an extension to the SSv2 dataset with more than $20K$ detailed annotations of core physical events in more than $10K$ videos. These annotations may be used in future efforts to develop new architectures or foundation models with cognitive understanding of visual dynamic interactions.
@@ 8.1 Action and object recognition
This section provides extended descriptions of the experimental setup of the evaluation on the tasks related to action and object recognition, mentioned in Section 4 of the main manuscript. All experiments were conducted using the extracted video frames of the full length videos from the original SSv2 dataset [10]. The models were given, as input, the sequence of video frames and prompted with the task-specific textual information detailed below. Five models including, GPT-4o, GPT-5.2, Gemini-2.5-Flash, Gemini-3-Pro and Qwen-2.5-VL-72B, were evaluated under the zero-shot regime (without any examples), and without any of the modifying conditions (Grounding and Reasoning).
@@
Listing LABEL:supplst:placeholder_prompt and LABEL:supplst:act_recog_prompt presents the prompts used in the experiments for testing the models performance in the tasks of object and action recognition, respectively. In the listed prompts, the strings template_sentence, template_id, and n_slots denote variables that were automatically replaced during execution for each experiment, according to the specific interaction template. The prompt structure itself was fixed, while these fields changed to reflect the corresponding template text, its numeric ID, and the number of placeholder slots.
@@ 8.2 Tow-Shot with Feedback
Inspired by recent work on boosting LLMs reasoning with verification feedback [32, 35, 12], we designed a variant to the common TS regime, in which the label of the second example was presented to the model indirectly, through an iterative feedback session. In this session, the model performed the detection task on the second example video and then prompted a numerical verification feedback by the user side. The feedback indicate a metric on the gap between the predicted frame ($f_{p}$) and the true frame ($f_{t}$) of the event. We defined the error in prediction, $\epsilon(f_{t},f_{p})$, via a sigmoid function, as shown in Eq. 1. The error function is shifted to the middle of the frame range ($\frac{N}{2}$) (as we constrained the error to be in the range of $[0,1]$, and required it to be $\epsilon(\frac{N}{2})=0.5$). The score was defined as shown in Eq. 2. The iterative feedback session ended when the model predicted the correct frame, or after it exceeded a limit of allowed trials $T_{th}$, which in our experiments was set to 10 (equivalent to the maximal number of iterations in the naive case where the model simply scans all the frames in turn until it gets to the right frame). After the iterative session ended, the model was instructed to perform the main detection task on the test video. In our experiment, we included examples only of other events from the same full video from which the test event was cropped, thus providing the model context with familiar context from the test video. The protocol followed the algorithm in Algorithm 1. The instructing prompt is presented in Listing LABEL:lst:ts_fb_prompt.
@@
Listing S1: Instruction prompt for object detection.
@@
Listing S2: Instruction prompt for event bounding-box detection.
@@
Listing S3: Instruction prompt for action recognition.
@@
8.3 Chain-of-Thought (CoT) Tuning
In the main experiment, examples included only videos with the true frame, but without further description of the dynamic event. In contrast, in this experiment we provided the agent with detailed steps for detecting the frame where the event occurs, for each example video. Taking inspiration from [32], in which the authors showed that an LMM can improve in the task of solving mathematical problems by presenting it examples of solutions to similar mathematical problems with a detailed, step-by-step solutions.
In our experiments we fed the agent a prompt that included three parts, viz. 1) a general explanation of the goal the agent was required to achieve, 2) a set of 1 — 8 examples with a full solution, and 3) the final test set, for which the final accuracy was computed (see LABEL:lst:cot-exp). For this matter we manually formulated 18 CoT prompts (9 for contact and 9 for release events), each with a detailed explanation of the scene and the interactions between a hand and objects located in it.
Listing S4: Instruction prompt for TS with feedback experiment.
User:
In the next procedure follow these rules:
6) You should never predict the same frame twice.
7) Stop your predictions only when the score is above 0.98.
Listing S5: The prompt template provided to the LMM agent in course of the CoT prompt tunning experiment. The prompt included an introduction describing a general goal the agent was required to achieve; a set of 1 to 8 examples, each including a detailed CoT, which described the objects in the scene and the steps towards the correct solution, and the corresponding frames of the examples; finaly, the test sequence was included as the last part of the prompt.
User:
You are an expert in video understanding and motion analysis.
We are interested in detecting and interpreting interactions in video sequences. We focus on interactions between hands and objects.
An interaction usually begins with a contact between a hand and a target object, when the static target object starts to move with a hand holding it.
An interaction usually ends when a hand releases a target object and starts to move without the object after it ungrasps the object.
You will be provided with consecutive frames from video sequences, where the order of the frames in each sequence is deteremined by the upload ordering and the acending index numbers in the file names.
You are asked to detect the exact frame where the interaction occurs. Your final answer should be brief in the format: "Prediction: <frame number>". Before your final answer, provide a step by step detailed description of all intermediate reasoning grounded in the video frames, in support of your final answer.
Example1:
<example video frame #1>
<example video frame #10>
STEPS:
1. Identify the Event:
< example text >
2. Frames analysis:
< example text >
3. Conclusion for prediction:
< example text >.
4. Final answer:
Prediction: < example true frame >
Example2:
Example8:
Test:
The following is the test video sequence.
<test video frame #1>
<test video frame #10>
Please detect to exact frame where the interaction occurs. Similar to the examples, provide a detailed reasoning before your final answer.
As before, at each iteration we requested the LMM agent to detect the exact frame in which the interaction (i.e., contact or release) happens, but this time, we provided it also with a subset of examples, with a varying length between 1 to 8, similarly to the experiments in [32]. In each iteration a random set of examples (of the same type as the test video) were chosen and their videos together with the corresponding CoT were sent as a prefix to the actual question. It is also worth to note that the examples were checked to be different from the test video in each iteration.
Text-based ablation.
LMMs strongly rely on the text context to answer questions. To establish a lower-bound for the LMMs’ expected performance, we conducted text-based ablation on blind model performance. We tried removing completely the images from the input prompt, as well as replacing the video images with blank images. However, in both cases the models detected the missing visual input and did not provide a prediction. The models’ response was something like: ”Actual images are not provided, therefore, this is a hypothetical analysis… Prediction: None”.